OpenStudy (anonymous):

A room contains 96 people. A man leaves and is replaced ny a women, leaving three times as many women as men in the room. How many men and women were there originally ?

OpenStudy (anonymous):

W:M=3:1=3*(96/4):(96/4)=72:24 Women: 71, Men: 25

Directrix (directrix):

Let M = the original number of Men Let W = the original number of Women ---------- M + W = 96 * 3( M - 1) = (W + 1) 3M - 3 = W + 1 3M - 4 = W ** -======= M + W = 96 * -- > W = 96 - M 3M - 4 = W ** Substitute 96 - M for W 3M - 4 = 96 - M 4M = 100 M = 25----->Orignal Number of Men W = 96 - M = 96 - 25 = 71 --> Original Number of Women